# On Ball’s conjectured Santaló type inequality

### v1

@@ 1 Introduction
For each $n\in\mathbb{N}$, we fix an inner product $\langle\cdot,\cdot\rangle$ and an orthonormal basis $\{e_{1},\dots,e_{n}\}$ in $\mathbb{R}^{n}$. The origin will be denoted by $o$. For $x\in\mathbb{R}^{n}$, set $x_{j}:=\langle x,e_{j}\rangle$, $j=1,\dots,n$, and $\|x\|:=\sqrt{\langle x,x\rangle}$. Set also $B_{2}^{n}:=\{x\in\mathbb{R}^{n}:\|x\|\leq 1\}$ and $S^{n-1}:=\{x\in\mathbb{R}^{n}:\|x\|=1\}$ to be the Euclidean unit ball and the Euclidean unit sphere, respectively, with respect to the inner product $\langle\cdot,\cdot\rangle$. Set $\mathbb{R}^{n}_{+}$ to be the first orthant, that is, $\mathbb{R}^{n}_{+}=\{x\in\mathbb{R}^{n}:x_{i}\geq 0,\ i=1,\dots,n\}$. A convex body $K$ will be a compact convex set with nonempty interior. In addition, $K$ will be called symmetric, if K=KK=-K. Finally, $K$ will be called unconditional if for any $x=(x_{1},\dots,x_{n})\in K$ and for any choice $\varepsilon_{1},\dots,\varepsilon_{n}\in\{\pm 1\}$, the point $(\varepsilon_{1}x_{1},\dots,\varepsilon_{n}x_{n})$ is also contained in $K$.
@@
To be more precise, (4) was originally established (as part of the proof of the unconditional case of Conjecture 1.1) in [5] and [6] for $X=C\cap\mathbb{R}^{n}_{+}$ and $Y=C^{\circ}\cap\mathbb{R}^{n}_{+}$, where $C$ is an unconditional convex body. However, the extension to the more general setting is straightforward (see e.g. Kalantzopoulos, Saroglou [17, Proposition 2.1] and the comments after it). We recall the argument for Theorem B, together with the characterization of the equality case in Section 2.
@@
Theorem 1.5 implies the stability of various additional linear invariant fundamental inequalities in terms of the symmetric difference metrics. Stability versions of the Blaschke-Santaló inequality and the Affine Isoperimetric inequality (cf. (12)) have been known in terms of the ”easier to handle” Banach-Mazur-distance (see Ball, Böröczky [7] and Böröczky [9]), but this is the first instance, following the footsteps of Fusco, Maggi, Pratelli [14] and Figalli, Maggi, Pratelli [13], to provide a stability version symmetric difference metrics, which is the most natural distance in this respect.
@@
$|K|\cdot|{\color[rgb]{0,0,0}K^{\circ}}|\leq\left(1-\theta_{n}A(K,E)^{2}\right)|B^{n}_{2}|^{2},$ (11)
@@
$\Omega(K)\leq$ $n|B^{n}_{2}|^{\frac{2}{n+1}}|K|^{\frac{n-1}{n+1}}$ (12)
@@
$\Omega(K)\leq$ $\left(1-\theta_{n}A(K,E)^{2}\right)n|B^{n}_{2}|^{\frac{2}{n+1}}|K|^{\frac{n-1}{n+1}}$ (13)
@@ 3 Yao-Yao partitions
is one to one. Following [19, Definition 6], we say (inductively) that a Yao-Yao partition $\cal P$} based on a hyperplane $F\subseteq\mathbb{R}^{n}$ is adapted to the coordinate system $(\lambda_{1},\dots,\lambda_{n})$ if $F=\{y\in E:\lambda_{1}(y)=\lambda_{1}(x)\}$ and $\cal P$$_{+}$, $\cal P$$_{-}$ are both adapted to $(\lambda_{2}|_{F},\dots,\lambda_{n}|_{F})$. According to Lehec’s [19, Proposition 14], if $\mu$ is a finite Borel measure on $\mathbb{R}^{n}$ such that $\mu(H)=0$ for all hyperplanes $H\subseteq\mathbb{R}^{n}$ and $\textnormal{supp}\,\mu=\mathbb{R}^{n}$, then there exists a Yao-Yao equipartition of $\mu$ that is adapted to $(\lambda_{1},\dots,\lambda_{n})$, and two such Yao-Yao partitions have the same center. In particular, the following crucial ingredient for the proof of Theorem 1.3 can be deduced from the results in [19].
@@
For an even and integrable function $f:\mathbb{R}^{n}\to(0,\infty)$ and for a positive integer $k$, let $\mu_{k}:=\mu+(1/k)f\,dx$, and hence $\mu_{k}$ is an even finite Borel measure that assigns zero mass to any hyperplane and $\textnormal{supp}\,\mu_{k}=\mathbb{R}^{n}$. According to [19, Proposition 14], there exists a unique Yao-Yao equipartition ${\cal P}_{k}$ of $\mu_{k}$ which is adapted $(\lambda_{1},\dots,\lambda_{n})$. Since $\mu_{k}$ is also even, uniqueness of the center ensures that for each $k$, the center of ${\cal P}_{k}$ is the origin (cf. [19, Lemma 15]). But then, approximation ([19, Lemma 12]) shows that $\mu$ also has a Yao-Yao equipartition $\cal P$} with center $o$, adapted to $(\lambda_{1},\dots,\lambda_{n})$. In particular, taking $\lambda_{1}(x)=\langle u,x\rangle$, $\cal P$} is based on the hyperplane
@@
We set $\omega_{n}=|B^{n}_{2}|$. We note that if $L\subseteq\mathbb{R}^{n}$ is a symmetric convex body in isotropic position, then Kannan, Lovász, Simonovits [18] prove that there exists an $o$-symmetric ball $B\subseteq L$ such that $L\subseteq nB$. If, in addition, $|L|=\omega_{n}$, then we deduce that
@@
Let $u\in S^{n-1}$ and $K\subseteq\mathbb{R}^{n}$ be a symmetric convex body, such that $K^{\circ}$ is in isotropic position with $|K^{\circ}|=|B^{n}_{2}|$, and let $\mathcal{P}$ be a Yao-Yao equipartition of the measure $d\mu(x)=\langle x,u\rangle^{2}\mathbf{1}_{K}(x)\,dx$ where $\mathcal{P}$ is centered at the origin, is based on $u^{\bot}$ and having axis $v\in S^{n-1}$. There exist explicit $C>0$, $R>n$ and $\varepsilon_{0}\in(0,1)$ depending only on $n$ satisfying that if
@@
As $K^{\circ}$ is in isotropic position with $|K^{\circ}|=|B^{n}_{2}|$, we deduce from (34) that
@@
There exists a positive definite symmetric matrix $\Phi\in{\rm GL}(n)$ such that $\Phi E^{\prime}=B^{n}$, and the eigenvalues of $\Phi$ are between $\frac{1}{R}$ and $R$, and hence $R^{-n}\leq\det\Phi\leq R^{n}$. In particular, the $o$-symmetric ellipsoid $\widetilde{E}=\Phi E$ and the convex cone $\tilde{\sigma}=\Phi\sigma$ satisfy
@@
and it is sufficient to find exlicit $\delta,C>0$ depending on $n,R,s$ such that
@@
as $e^{t}\leq 1+2t$ for $t\in(0,1)$, verifying (63) with $C=\frac{3nR^{n}|B^{n}_{2}|}{c}$, and in turn proving Lemma 6.2. ∎
@@
We may assume that $K^{\circ}$ is in isotropic position and $|K^{\circ}|=|B^{n}_{2}|$, and hence (34) yields that
@@
$|K|\cdot|K^{\circ}|\leq\left(1-\theta_{n}\cdot\widetilde{A}(K,B^{n}_{2})^{2}\right)|B^{n}_{2}|^{2}$ (102)
@@
(iii) $|K_{t}\Delta E|\geq c_{0}|t|$ for any $o$-symmetric ellipsoid $E\subseteq\mathbb{R}^{n}$ with $|E|=|B^{n}_{2}|$;
(iv) $|K_{t}^{\circ}|\geq(1-C_{0}t^{2})|B^{n}_{2}|$.
@@
$|K_{t}|\cdot|K_{t}^{\circ}|\geq\left(1-C\cdot\widetilde{A}(K_{t},B^{n}_{2})^{2}\right)|B^{n}_{2}|^{2}\mbox{ \ for \ }C=\frac{C_{0}}{c_{0}|B^{n}_{2}|^{1/n}},$
@@
$K_{t}=\lambda_{t}\cdot\widetilde{K}_{t}\mbox{ \ for }\lambda_{t}=\frac{|B^{n}_{2}|^{\frac{1}{n}}}{|\widetilde{K}_{t}|^{\frac{1}{n}}}\in\left(\frac{3}{4},\frac{4}{3}\right)$
satisfies that $|K_{t}|=|B^{n}_{2}|$, $\frac{1}{2}B^{n}_{2}\subseteq K_{t}\subseteq 2B^{n}_{2}$; moreover, $\lambda_{t}$ and $|K_{t}^{*}|$ are $C^{\infty}$ functions of $t\in(-\eta,\eta)$, and hence $\lambda_{t}=1+O(|t|)$, and
@@
Finally, we prove (iii) indirectly; therefore, we suppose that there exists a sequence $t_{k}\neq 0$ tending to zero as $k$ tends to infinity and a sequence of $o$-symmetric ellipsoids $E_{t_{k}}$ such that $|E_{t_{k}}|=|B^{n}_{2}|$ and
@@
and hence the triangle inequality for the symmetric difference metric and $\lambda_{t_{k}}\in(\frac{3}{4},\frac{4}{3})$ imply that $\big|B^{n}_{2}\Delta\left[(1+t_{k})B^{n}_{2}\right]\Big|=o(|t_{k}|)$. However, $\big|B^{n}_{2}\Delta\left[(1+t_{k})B^{n}_{2}\right]\Big|\geq|t_{k}|\cdot|B^{n}_{2}|$, and this contradiction finally verifies (iii).

### v2

@@ 1 Introduction
For each $n\in\mathbb{N}$, we fix an inner product $\langle\cdot,\cdot\rangle$ and an orthonormal basis $\{e_{1},\dots,e_{n}\}$ in $\mathbb{R}^{n}$. The origin will be denoted by $o$. For $x\in\mathbb{R}^{n}$, set $x_{j}:=\langle x,e_{j}\rangle$, $j=1,\dots,n$, and $\|x\|:=\sqrt{\langle x,x\rangle}$. Set also $B_{2}^{n}:=\{x\in\mathbb{R}^{n}:\|x\|\leq 1\}$ and $S^{n-1}:=\{x\in\mathbb{R}^{n}:\|x\|=1\}$ to be the Euclidean unit ball and the Euclidean unit sphere, respectively, with respect to the inner product $\langle\cdot,\cdot\rangle$. and denote $\omega_{n}=|B^{n}_{2}|$. Set $\mathbb{R}^{n}_{+}$ to be the first orthant, that is, $\mathbb{R}^{n}_{+}=\{x\in\mathbb{R}^{n}:x_{i}\geq 0,\ i=1,\dots,n\}$. A convex body $K$ will be a compact convex set with nonempty interior. In addition, $K$ will be called symmetric, if K=KK=-K. Finally, $K$ will be called unconditional if for any $x=(x_{1},\dots,x_{n})\in K$ and for any choice $\varepsilon_{1},\dots,\varepsilon_{n}\in\{\pm 1\}$, the point $(\varepsilon_{1}x_{1},\dots,\varepsilon_{n}x_{n})$ is also contained in $K$.
@@
To be more precise, (4) was originally established (as part of the proof of the unconditional case of Conjecture 1.1) in [5] and [6] for $X=C\cap\mathbb{R}^{n}_{+}$ and $Y=C^{\circ}\cap\mathbb{R}^{n}_{+}$, where $C$ is an unconditional convex body. However, the extension to the more general setting is straightforward (see e.g. Kalantzopoulos, Saroglou [18, Proposition 2.1] and the comments after it). We recall the argument for Theorem B, together with the characterization of the equality case in Section 2.
@@
Theorem 1.5 implies the stability of various additional linear invariant fundamental inequalities in terms of the symmetric difference metrics. Stability versions of the Blaschke-Santaló inequality and the Affine Isoperimetric inequality (cf. (12)) have been known in terms of the ”easier to handle” Banach-Mazur-distance (see Ball, Böröczky [7], Böröczky [9] and Ivaki [17]), but this is the first instance, following the footsteps of Fusco, Maggi, Pratelli [14] and Figalli, Maggi, Pratelli [13], to provide a stability version symmetric difference metrics, which is the most natural distance in this respect.
@@
$|K|\cdot|{\color[rgb]{0,0,0}K^{\circ}}|\leq\left(1-\theta_{n}A(K,E)^{2}\right)\omega_{n}^{2},$ (11)
@@
$\Omega(K)\leq$ $n\omega_{n}^{\frac{2}{n+1}}|K|^{\frac{n-1}{n+1}}$ (12)
@@
$\Omega(K)\leq$ $\left(1-\theta_{n}A(K,E)^{2}\right)n\omega_{n}^{\frac{2}{n+1}}|K|^{\frac{n-1}{n+1}}$ (13)
@@ 3 Yao-Yao partitions
is one to one. Following [20, Definition 6], we say (inductively) that a Yao-Yao partition $\cal P$} based on a hyperplane $F\subseteq\mathbb{R}^{n}$ is adapted to the coordinate system $(\lambda_{1},\dots,\lambda_{n})$ if $F=\{y\in E:\lambda_{1}(y)=\lambda_{1}(x)\}$ and $\cal P$$_{+}$, $\cal P$$_{-}$ are both adapted to $(\lambda_{2}|_{F},\dots,\lambda_{n}|_{F})$. According to Lehec’s [20, Proposition 14], if $\mu$ is a finite Borel measure on $\mathbb{R}^{n}$ such that $\mu(H)=0$ for all hyperplanes $H\subseteq\mathbb{R}^{n}$ and $\textnormal{supp}\,\mu=\mathbb{R}^{n}$, then there exists a Yao-Yao equipartition of $\mu$ that is adapted to $(\lambda_{1},\dots,\lambda_{n})$, and two such Yao-Yao partitions have the same center. In particular, the following crucial ingredient for the proof of Theorem 1.3 can be deduced from the results in [20].
@@
For an even and integrable function $f:\mathbb{R}^{n}\to(0,\infty)$ and for a positive integer $k$, let $\mu_{k}:=\mu+(1/k)f\,dx$, and hence $\mu_{k}$ is an even finite Borel measure that assigns zero mass to any hyperplane and $\textnormal{supp}\,\mu_{k}=\mathbb{R}^{n}$. According to [20, Proposition 14], there exists a unique Yao-Yao equipartition ${\cal P}_{k}$ of $\mu_{k}$ which is adapted $(\lambda_{1},\dots,\lambda_{n})$. Since $\mu_{k}$ is also even, uniqueness of the center ensures that for each $k$, the center of ${\cal P}_{k}$ is the origin (cf. [20, Lemma 15]). But then, approximation ([20, Lemma 12]) shows that $\mu$ also has a Yao-Yao equipartition $\cal P$} with center $o$, adapted to $(\lambda_{1},\dots,\lambda_{n})$. In particular, taking $\lambda_{1}(x)=\langle u,x\rangle$, $\cal P$} is based on the hyperplane
@@
We note that if $L\subseteq\mathbb{R}^{n}$ is a symmetric convex body in isotropic position, then Kannan, Lovász, Simonovits [19] prove that there exists an $o$-symmetric ball $B\subseteq L$ such that $L\subseteq nB$. If, in addition, $|L|=\omega_{n}$, then we deduce that
@@
Let $u\in S^{n-1}$ and $K\subseteq\mathbb{R}^{n}$ be a symmetric convex body, such that $K^{\circ}$ is in isotropic position with $|K^{\circ}|=\omega_{n}$, and let $\mathcal{P}$ be a Yao-Yao equipartition of the measure $d\mu(x)=\langle x,u\rangle^{2}\mathbf{1}_{K}(x)\,dx$ where $\mathcal{P}$ is centered at the origin, is based on $u^{\bot}$ and having axis $v\in S^{n-1}$. There exist explicit $C>0$, $R>n$ and $\varepsilon_{0}\in(0,1)$ depending only on $n$ satisfying that if
@@
As $K^{\circ}$ is in isotropic position with $|K^{\circ}|=\omega_{n}$, we deduce from (34) that
@@
There exists a positive definite symmetric matrix $\Phi\in{\rm GL}(n)$ such that $\Phi E^{\prime}=B^{n}_{2}$, and the eigenvalues of $\Phi$ are between $\frac{1}{R}$ and $R$, and hence $R^{-n}\leq\det\Phi\leq R^{n}$. In particular, the $o$-symmetric ellipsoid $\widetilde{E}=\Phi E$ and the convex cone $\tilde{\sigma}=\Phi\sigma$ satisfy
@@
and it is sufficient to find explicit $\delta,C>0$ depending on $n,R,s$ such that
@@
as $e^{t}\leq 1+2t$ for $t\in(0,1)$, verifying (63) with $C=\frac{3nR^{n}\omega_{n}}{c}$, and in turn proving Lemma 6.2. ∎
@@
We may assume that $K^{\circ}$ is in isotropic position and $|K^{\circ}|=\omega_{n}$, and hence (34) yields that
@@
$|K|\cdot|K^{\circ}|\leq\left(1-\theta_{n}\cdot\widetilde{A}(K,B^{n}_{2})^{2}\right)\omega_{n}^{2}$ (102)
@@
(iii) $|K_{t}\Delta E|\geq c_{0}|t|$ for any $o$-symmetric ellipsoid $E\subseteq\mathbb{R}^{n}$ with $|E|=\omega_{n}$;
(iv) $|K_{t}^{\circ}|\geq(1-C_{0}t^{2})\omega_{n}$.
@@
$|K_{t}|\cdot|K_{t}^{\circ}|\geq\left(1-C\cdot\widetilde{A}(K_{t},B^{n}_{2})^{2}\right)\omega_{n}^{2}\mbox{ \ for \ }C=\frac{C_{0}}{c_{0}\omega_{n}^{1/n}},$
@@
$K_{t}=\lambda_{t}\cdot\widetilde{K}_{t}\mbox{ \ for }\lambda_{t}=\frac{\omega_{n}^{\frac{1}{n}}}{|\widetilde{K}_{t}|^{\frac{1}{n}}}\in\left(\frac{3}{4},\frac{4}{3}\right)$
satisfies that $|K_{t}|=\omega_{n}$, $\frac{1}{2}B^{n}_{2}\subseteq K_{t}\subseteq 2B^{n}_{2}$; moreover, $\lambda_{t}$ and $|K_{t}^{*}|$ are $C^{\infty}$ functions of $t\in(-\eta,\eta)$, and hence $\lambda_{t}=1+O(|t|)$, and
@@
Finally, we prove (iii) indirectly; therefore, we suppose that there exists a sequence $t_{k}\neq 0$ tending to zero as $k$ tends to infinity and a sequence of $o$-symmetric ellipsoids $E_{t_{k}}$ such that $|E_{t_{k}}|=\omega_{n}$ and
@@
and hence the triangle inequality for the symmetric difference metric and $\lambda_{t_{k}}\in(\frac{3}{4},\frac{4}{3})$ imply that $\big|B^{n}_{2}\Delta\left[(1+t_{k})B^{n}_{2}\right]\Big|=o(|t_{k}|)$. However, $\big|B^{n}_{2}\Delta\left[(1+t_{k})B^{n}_{2}\right]\Big|\geq|t_{k}|\cdot\omega_{n}$, and this contradiction finally verifies (iii).
Acknowledgement: Böröczky’s research is supported in part by NKFIH NKKP grant 150613.